Monitoring atmospheric composition & climate
 
 
Change log for fnyp

The MACC NRT o-suite (experiment fnyp) was started on 2012-03-21 for the analysis cycle 20111201. It was officially implemented as o-suite on 2012-07-05. On 7 October 2013 a new cycle (CY38R2) was introduced. On 24 February 2014 CY40R1 was introduced. Main features are:

  • runs in near-real-time
  • 120 h long forecast from 00UTC
  • IFS cycle 40R1
  • Coupled to Mozart V3.5
  • MACCity anthropogenic emissions - scaled over North America and Europe
  • GFASv1.0 fire emissions
  • IFS model resolution T255L60
  • CTM resolution 1.125 degrees x 1.125 degrees, L60
  • analysis resolution T159/T95, using 12h 4D-Var and VarBC
  • time period 20111201 -
  • Blacklist: black_ds_macc_nrt_40r1
  • branch: stj_CY37R3_MACC_NRT_aer_CO; stj_CY38R2_MACC_esuite; stj_CY40R1_MACC_esuite

 

Satellite data usage for atmospheric composition

Aerosol
Assimilated: AQUA MODIS, TERRA MODIS
   
Ozone
Assimilated: SBUV/2 21-layer, OMI, MLS, GOME-2A, GOME-2B (since 20140512)
Monitored: SEVIRI
   
Carbon Monoxide
Assimilated: IASI-A, MOPITT
Monitored: IASI-B
   
NO2
Assimilated: OMI
Monitored: GOME-2A, GOME-2B
   
SO2
Assimilated: OMI
Monitored: GOME-2A, GOME-2B
   
HCHO  
Monitored: GOME-2A

 

Changes introduced during production affecting the quality of the products

real date exp date description of change
20120321 2011120100 Start of 37R3 e-suite
20120321 20120221 MLS instrument offline after suspected Single Event Upset event that shut down the on-board computer at 18:39UTC on Sunday 19th February 2012. The current plan is for the instrument to return to normal science operations around Friday February 24th 2012. No data will be available during this period.
20120321 20120315 NO MLS data. AURA MLS Instrument has encountered an anomaly as of DOY 074 1834z and MLS has been off since 14/03. Due back on 21/03 according to JPL website
20120321 20120305 No MOPITT CO data because of hot calibration event between 5-16 March.
20120322 2012032200 MLS back. MLS Instrument has been stored to "Nominal" scan mode and it is now back on-line.
MLS Level 2 NRT Data are available again starting from 2012-03-21 18:10.
20120330 20120330 NO IASI CO data between ~14:00 - 21.00 because of problems with LATMOS data server.
20120321 2011120100 Replacement of flv2 with fnyp. Purpose of change was to have improved aerosol model and scaled CO emissions over North America and Europe.
20120416 20120408 No Envisat data (e.g. SCIAMACHY, MIPAS) since 8 April after contact with satellite was lost.
20120416 20120423 No IASI CO data. The IASI CO NRT processing will stop Monday, 23 April at 7 UTC because of a power outage. If all goes well, this will only last 48 to 72 hours and data will come back on Wednesday, 25 April.
20120423 2012031500 Switched off verify jobs, because there is still a problem with retrieving Z from MARS when NOx is there as well.
20120606 2012060600 OH added to list of output variables of MOZART forecast
20120608 2012060800 Updates to aerosol model to optimize sources as well as sedimentation, wet and dry deposition
20120612

20120524-

20120612

Reduced MOPITT CO data numbers between 20120524-20120612 because bug in MOPITT processing chain led to many (but not all) files being generated with no data inside.
20120612

20120601 -

20120611

No MLS data. Tightening of security on NASA's EOSDIS User Registration System required a password change on 1 June 2012. By mistake this was not applied to the acquisition of MLS ozone until 11 June, and not data were picked up during this period.
20120627 20120627-

LANCE-MLS Registered Users: Please expect gaps in the MLS Level 2 NRT data. There has been a resurgence in activity by the existing anomaly on the spacecraft, that Aura has been experiencing for the past couple years. GES DISC team continues to investigate the cause(s) of the gaps to recover as much NRT data as possible.

20120701 20120702

Planned OMI data outage between July 02, 2012 , 1749 UTC and
July 03, 2012, 1749 UTC

20120704 20120703

Bugfix for IASI CO observation errors. The errors had been read as relative error in % and set to max(5%, obserr)  in the BUFR software. They are however given as relative error in the original data files. This meant that prior to today all errors were a factor 100 too small and set to 5%. First observation with proper relative error is for 20120703 13:17:57.

20120705 2012070500 Introduced 37R3 as new MACC o-suite
20120709 2012070900 Bugfix in model script for UV output
20120725 July 2012

Reduced MOPITT data volume because of several server upgrades by the MOPITT team during July. These  have caused some data outages and a lot of data arrived later than usual and did not make the extraction cut-off time.

20120802 2012080200 Technical bugfix to correct NOx background standard deviations being overwritten by O3 standard deviations. Impact analysis from short parallel run showed small but significant change in upper troposphere and lower stratosphere in the tropics. The bug also affected CO, but with only minor impact.
20121002 2012100200 Recompilation of MOZART-3.5 to fix bug in HNO3 output.
20121003 2012100200

Failure in pregeos because of new GOES-14 data. Added GOES 14 to script and blacklist. Files rtcoef_goes_14_imager.dat and sccldcoef_goes_14_imager.dat added to ~rdx/data/37r3/sat/rttov and mklinks rerun.

20130107 2013010706

New MLS V3.4 data. First V3.4 MLS data are for 20130107, 03:11:11. The new version has more vertical levels and useful data down to 261 hPa (instead of 68 hPa). This should improve the vertical structure of the ozone analysis.

20130129 2013012900 Move fnyp to new HPCF system
20130129 2013012900

New MOPITT v5 CO data, because production of MOPITT V4 data will stop soon. Initial tests showed MOPITT V5 data to have a smaller bias relative to the model (and to IASI data). More details about the V5 data can be found in www.acd.ucar.edu/mopitt/v5_users_guide_beta.pdf .

20130312 2013031200 Begin of passive monitoring of GOME-2 HCHO.
20130818 2013081800 Remove MOPITT CO data in fetchobs because bug in the observation operator (that is fixed in 38r2) leads to failures. Removing the data was the simplest solution while the DA people were on holiday. No MOPITT CO data used after 2013081800.
20131003 20131003 Blacklist MLS O3 because Aura LANCE is experiencing data-related processing failures and the number of data received is greatly reduced.
20131007 2013100700 Introduction of CY38R2 from e-suite fwu0 based on branch stj_CY38R2_MACC_esuite
20131210 2013121000 Begin of passive monitoring of GOME-2 MetOP-B O3.
20131217 2013121700 Add METOP-B GOME-2 to obstat monitoring (change of general.cfg file)
20140224 2014022400 Introduction of CY40R1 from e-suite fzpr based on branch stj_CY40R1_MACC_esuite
20140225 2014022400 Fixed problem with missing multiple wavelength AOD
20140319 2014031900 Correction of bug in ozone increments from IASI/AIRS radiances
20140512 2014051200 Activate Metop-B GOME-2 O3
20140605 2014060412 Failure in screening with U-wind observations. Looks like a technical failure; increased NLLNDA from 1000 to 1200
20140904 2014090100 - 2014090400 From 29 August 2014 onwards the MODIS FRP algorithm interpreted the hot lava on Iceland as a wildfire. Consequently, the MACC-III GFAS fire emission analysis system started producing wildfire emissions from this date onwards. The MACC-III NRT analyses and forecasts were affected by this error from 1 September onwards. A fix was put in place for the GFAS analysis of 3 September. This meant that the NRT analyses and forecasts did not have these erroneous emissions anymore from 4 September onwards. However, some of the previously produced plumes will linger in the analyses due ot the transfer of information through a short forecast from one analysis cycle to the next.

 

Blacklist changes

real date
  expdate
description of change
20111111 20111101 Created black_ds_macc_nrt_37r3 from operational blacklist. Blacklist directory ec:/std/blacklist
20111206 20111206 Activate the assimilation of MIPAS O3 data. First cycle with active MIPAS data: 20111206, 12z.
20111206 20111206 Blacklist CSR/AMV from GOES-11, un-blacklist AMSU-A NOAA-15, blacklist MHS on METOP-A
20111208 20111208 Reactivate MHS on METOP-A
20111214 20111214 Blacklist HIRS on NOAA-17 and some METOP-A sensors
20111223 20111223 blacklisting of MTSAT-1R data
20120110 20120110

Reactivate MTSAT-2

20120119 20120119 Blacklist change  for GOES-15
20120124 20120124 Blacklist AMSU-A on NOAA-15 (gyro test)
20120130 20120130 Re-activate AMSU-A on NOAA-15
20120208 20120208 reintroduction of ground based GPS in passive mode
20120221 20120220 Blacklist NOAA-18 and METEOSAT-9 data until 20120227 because of  enhanced GYE Flight Software On-Orbit Test. This also affects SBUV/2 ozone data from NOAA-18.
20120426 20120425 Activate GOES-15 AMV, CSR.
20120426 20120426 Blacklist HIRS on NOAA-19 due to increased noise.
20120503 20120503 Reactivate HIRS on NOAA-19 because noise went down again.
20120625 20120625 Blacklist MET-9 CSR, overcast and AMVs from 25th June due to MET-9 north south orbit inclination manoeuvre.
20120717 20120717 Changed blacklist for SO2 from OMI so that the data are blacklisted for SOE<20 (was SOE<15 before, but that let in 1 dodgy observation over S-Pacific on 20120714, 12z.)
20121217 20121217 Blacklist SBUV/2 from NOAA-18 because the instrument has been affected by a mechanical anomaly that prevents any measurements. This action is taken as a precaution for when dissemination will restart.
20130128 20130128 Blacklist SBUV/2 from NOAA-17 because of restricted data coverage and much reduced number of good observations.
20131210 20131210 Include METOP-B GOME-2 O3 for monitoring
20140122 20140122 Blacklist GOME-2A O3 after GOME-2A switched itself off after a radiation triggered single event.
20140129 20140129 Re-activate GOME-2A O3.
20140325 20140325 The MOPITT instrument team will be performing the annual decontamination and calibration operations on the instrument from March 24 through April 8, inclusive. There will probably be no science products released for this time period. MOPITT CO blacklisted until we are sure there are no problems with the data when they come back.
20140429 20140429 Re-activate MOPITT CO.
20140512 20140512 Activate GOME-2B O3.
20140604 20140604 Blacklist night time MOPITT CO over Greenland. Problems with the cloud detection over polar surface at night can lead to high night time CO values over polar surfaces, esp Greenland.